OpenStudy (anonymous):

Find all solutions for the absolute value equation: I 2x-3I=x+3 a. {0} b. {0, -6} c. {0, 6} d. {6}

OpenStudy (hoblos):

I 2x-3I=x+3 so 2x-3=x+3 or 2x-3=-(x+3) x=6 2x-3=-x-3 3x = 0 x=0 thus solutions are {0, 6}